# Mitigating Visual Knowledge Forgetting in MLLM Instruction-tuning via Modality-decoupled Gradient Descent

### v1

@@ 5.2. Regularized Gradient Descent
(10) $\pi_{\theta}\leftarrow\pi_{\theta}-\nabla_{\theta}\mathcal{L}_{vl}(\theta)-\nabla_{\theta}\mathcal{L}_{v}(\theta)-\tilde{g}_{\theta}.$
5.3. Enabling Parameter-efficient Fine-tuning of MDGD via Gradient Masking
Parameter-efficient fine-tuning (PEFT) methods, such as adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021), aim to reduce the computational cost and memory usage when fine-tuning models on downstream tasks under practical constraints (Han et al., 2024). However, due to the requirement of directly estimating gradient directions on the pre-trained model parameters, MDGD cannot be directly applied to these PEFT methods, which introduce additional model parameters whose gradients are separate from the original model weights.
To address this challenge, we propose a variant, MDGD-GM, by formulating the gradient regularization term in Eq. (9) as gradient masking that selects model weights with efficient gradient directions. Specifically, we define the gradient mask as
(11) $M_{\tilde{g}_{\theta}}=\mathbf{1}\left\{\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|\|\bar{g}_{\theta}\|}\geq T_{\alpha}\right\},$
@@ 5.3. Enabling Parameter-efficient Fine-tuning of MDGD via Gradient Masking
(12) $\pi_{\theta}\leftarrow\pi_{\theta}-M_{\tilde{g}_{\theta}}\cdot\left(\nabla_{\theta}\mathcal{L}_{vl}(\theta)+\nabla_{\theta}\mathcal{L}_{v}(\theta)\right).$
@@
Table 1. Performance on various pre-trained tasks of LLaVA-1.5 models fine-tuned on Flickr30K and OKVQA. We report the best performance for each task in a bold font while the second best performance underlined.
6. Experiments
In this section, we conduct experiments on various datasets and backbone MLLMs to investigate the following research questions:
  1. (1)
    RQ1 (Overall Performance) Can MDGD prevent visual forgetting while improving downstream tasks?
  2. (2)
    RQ2 (Ablation Study) How do the visual alignment and gradient masking affect the MDGD’s performance?
  3. (3)
    RQ3 (Representation Learning) How does MDGD benefit effective multimodal representation learning in MLLMs?
  4. (4)
    RQ4 (Sensitivity Study) How does gradient masking ratio $\alpha$ affects the learning of MDGD?
  • Pre-trained Tasks: VQAv2 (Goyal et al., 2017), GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), SQA (Lu et al., 2022), TextVQA (Singh et al., 2019), POPE (Li et al., 2023a), and MM-Bench (Liu et al., 2023).
  • Pre-trained Tasks: VizWiz, OKVQA, A-OKVQA (Schwenk et al., 2022), Text-VQA, IconQA (Lu et al., 2021), POPE, and MM-Bench.
Baselines We compare our approach against several baselines:
  • Standard Fine-Tuning. For a fair comparison, we follow the setting of Model-Tailor (Zhu et al., 2024), where LLaVA-1.5 is fine-tuned on the last 6 layers and its feature adapter, with a total of 1.2B parameters. MiniCPM is fine-tuned on the last 8 layers and its feature resampler, with 517M parameters.
6.1. Overall Performance (RQ1)
6.1.1. Larger MLLM adapts better to downstream tasks but is more prone to visual forgetting.
@@ 6.1. Overall Performance (RQ1)
We study the visual forgetting problem on the LLaVA-1.5 MLLM which contains 7B model parameters and report performance comparison results in Table 1. We observe that the pre-trained LLaVA enables efficient instruction tuning on target tasks, where the zero-shot performance is near zero. When the model is fine-tuned on the image caption task, Flickr30K, which largely differs from the pre-trained tasks of visual question-answering, the model can learn a degraded multimodal representation, which causes visual forgetting in its projected visual representation space (in Section 4) and its average performance on pre-trained tasks drops 33.63% compared with zero-shot performance. Fine-tuning on visual question-answering task OKVQA, which is similar to the pre-trained tasks, can also cause a 13.44% performance drop, due to the limited image-text pairs existing in the downstream task, which potentially leads to MLLM’s visual understanding drift.
6.1.2. Smaller MLLM also experiences visual forgetting while limited in downstream task improvements.
To validate the observation on a smaller MLLM, we report the comparison results of MiniCPM-V-2.0 with 2.8B model parameters in Table 2. We observe that compared with the LLaVA MLLM, MiniCPM suffers from less prominent visual forgetting. The average performance drop of the model limits to 6.28% and 4.25% when fine-tuning on PathVQA and TextCaps, respectively. We attribute this observation to MiniCPM learning a more compact and constrained visual representation space during pre-training, causing the visual representations of target task images to be less aligned with those of the pre-trained MLLM. Consequently, MiniCPM exhibits limited improvement in downstream tasks, as its restricted ability to acquire additional visual knowledge leads to ineffective instruction tuning.
6.1.3. MDGD prevents visual forgetting while maintaining downstream task improvements.
By employing MDGD in MLLM instruction tuning, we observe the LLaVA’s average performance drop on pre-trained tasks reduces to 3.59% when fine-tuned on OKVQA and also achieves a 1.45% improvement when fine-tuned on Flickr30K, which demonstrates the efficiency of MDGD in mitigating visual forgetting. For the smaller MLLM, MiniCPM, MDGD achieves comparable downstream task improvements with direct fine-tuning, while completely eliminating visual forgetting in the pre-trained tasks. MDGD and its variants consistently achieve the best average performance for both MLLMs, demonstrating its great potential for incremental learning on individual downstream tasks.
6.1.4. Comparison with baseline methods.
In Table 1, we compare MDGD with LoRA fine-tuning and Model Tailor (Zhu et al., 2024) on LLaVA-1.5, which are designed for parameter-efficient fine-tuning. We observe that LoRA fine-tuning can suffer from significant visual forgetting on Flickr30K and OKVQA. Since LoRA introduces additional representation projections in intermediate layers, the pre-trained multimodal representations can be projected into a lower-rank subspace leading to visual forgetting (in Section 4), due to the limitation of image-text pairs in the target dataset. Model Tailor is designed for MLLM anti-forgetting, which identifies “patches” of sub-model parameters significantly affected by fine-tuning on the target task. However, since the method is not specifically designed for MLLMs, the unique challenge of visual forgetting cannot be effectively mitigated while maintaining robust performance on the target task.Thus, we observe that Model Tailor’s performance is sensitive to the target task datasets (e.g., better on Flickr30K than OKVQA), whereas MDGD consistently outperforms Model Tailor in terms of both average task scores and H-scores across the two datasets. In Table 2, we also report the results of the MDGD comparison with LoRA fine-tuning on MiniCPM. We observe consistent improvements on the average task performance of MDGD when fine-tuned on PathVQA and TextCaps, especially MDGD achieves 2.43% and 1.83% on the pre-trained tasks of PathVQA and TextCaps, respectively, which demonstrates the effectiveness of MDGD in mitigating visual forgetting.
@@
Table 2. Performance on various pre-trained tasks of MiniCPM-V2.5 models fine-tuned on PathVQA and TextCaps. We report the best performance for each task in a bold font while the second best performance underlined.
@@
Figure 3. T-SNE plots of the distribution of extracted visual $\pi(X^{v})$ and multimodal $z^{vl}$ representations from pre-trained LLaVA-1.5, and models with direct fine-tuning and MDGD on OKVQA and Flickr30K.
(a) $z^{vl}$ on PathVQA
(b) $\pi(X^{v})$ on PathVQA
(c) $z^{vl}$ on TextCaps
(d) $\pi(X^{v})$ on TextCaps
Figure 4. T-SNE plots of the distribution of extracted visual $\pi(X^{v})$ and multimodal $z^{vl}$ representations from pre-trained MiniCPM, and models with direct fine-tuning and MDGD on PathVQA and TextCaps.
Figure 5. Illustration of (a) the learning process of three methods based on task loss $\mathcal{L}_{vl}(\phi,\theta)$, (b) the average regularized cosine similarity $\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|\|\bar{g}_{\theta}\|}$ in Eq.(11) for gradient masking at varying ratios, and (c) the visual representation loss $\mathcal{L}_{v}(\phi,\theta)$ in Eq.(6) for gradient masking at varying ratios $\alpha$.
(a) LLaVA models pretrained, finetuned, and fine-tuned with MDGD
(b) MiniCPM models pretrained, finetuned, and fine-tuned with MDGD
Figure 6. The effective rank comparison on individual downstream fine-tuning datasets.
6.2. Ablation Study (RQ2)
6.2.1. Ablation study on visual alignment.
We compare MDGD with its two variants, MDGD w/o visual align and MDGD-GM. MDGD w/o visual align enables MDGD without including visual representation loss $\mathcal{L}_{v}(\phi,\theta)$ Eq.(6), to understand the effect of directly optimizing to reduce the visual representation discrepancy between the current model and pre-trained model. We observe that MDGD w/o visual align maintains relatively comparable performance to MDGD on OKVQA and PathQA, due to the reduced need for visual representation adaptation in such visual question-answering tasks. In contrast, tasks like image captioning on Flickr30K and TextCaps benefit from feature alignment regularization, which directly mitigates visual understanding drift in the MLLM.
6.2.2. Ablation study on gradient masking.
The other variant, MDGD-GM, leverages gradient masking to enable parameter-efficient fine-tuning (PEFT). We observe the PEFT variant of MDGD consistently achieves comparable performance across all tasks and backbone MLLMs, which only fine-tunes a subset of 10% original MLLM parameters used for direct fine-tuning and original MDGD. Different from conventional PEFT methods such as adapters, MDGD and its variants do not introduce additional parameters to the original model architecture, enabling continuous and incremental learning in an online setting (Maltoni and Lomonaco, 2019; Gao et al., 2023).
6.3. Representation Study (RQ3)
6.3.1. T-SNE Analysis on Visual Representation
To analyze the learning of visual and multimodal representation distributions in MLLMs, we create T-SNE (Van der Maaten and Hinton, 2008) plots to visualize the feature distributions extracted from pre-trained MLLMs, as well as MLLMs after standard fine-tuning and MDGD We illustrate the distributions of the multimodal features $z^{vl}$ extracted from the last token of the multimodal instruction tokens, and the visual features $\pi_{\theta}(X^{v})$ extracted from the last token of the input image tokens. We observe a consistent visual understanding drift in the MLLMs’ visual representation spaces after standard fine-tuning on Flickr30K and OKVQA with LLaVA (Figure 3b and 3d), as well as PathVQA and TextCaps with MiniCPM (Figure 4b and 4d). By employing MDGD to mitigate visual forgetting, we observe that visual understanding drift is effectively reduced, allowing the fine-tuned MLLM to retain pre-trained visual capabilities and preserve visual information.
We further observe a distributional discrepancy in the multimodal representation $z^{vl}$ of LLaVA (Figures 3a and 3c) between MDGD and the pre-trained MLLM. This discrepancy arises from the alignment of the MLLM to the target task through multimodal instructions, demonstrating effective adaptation to the downstream task of the LLaVA model. In addition, we also observe such multimodal distribution discrepancy reduces in a smaller MLLM, MiniCPM. This observation aligns with our findings on MiniCPM in Section 6.1, where we noted limited effects in model adaptation to downstream tasks. However, applying MDGD to MiniCPM mitigates visual forgetting by preventing degradation of both image and multimodal encodings into lower-rank representation spaces.
6.3.2. Effective Rank Analysis on Visual Representation
To quantitatively analyze the visual forgetting problem (in Section 4) described in Eq. (5), we calculate effective ranks of the visual representations extracted from the last hidden layer on the position of image tokens in individual MLLMs. We show the comparison results of LLaVA models in Figure 6(a) and MiniCPM models in Figure 6(b). We observe that with both the backbone models of LLaVA and MiniCPM, directly fine-tuning the pre-trained models on downstream tasks can lead to a consistent reduction of effective ranks in visual representations. Such observations validate the hypothesis in Section 4 regarding the potential visual forgetting problem in MLLM instruction tuning. In addition, we can observe that MDGD achieves consistent improvements in effective ranks compared with the standard fine-tuning method for both backbone MLLMs across various pre-trained tasks. In Figure 6(a), we observe that MDGD achieves comparable or even better effective ranks on pre-trained tasks, compared with the pre-trained LLaVA model. However, MDGD on MiniCPM in Figure 6(b) also suffers from the visual representation degradation problem, while MDGD consistently alleviates the problem. Such observation suggests a higher risk of visual forgetting in smaller-scale MLLMs.
6.4. Sensitivity Study (RQ4)
7. Conclusion

### v2

@@ 3.2 Regularized Gradient Descent
$\pi_{\theta}\leftarrow\pi_{\theta}-\nabla_{\theta}\mathcal{L}_{vl}(\theta)-\nabla_{\theta}\mathcal{L}_{v}(\theta)-\tilde{g}_{\theta}.$ (9)
3.3 Enabling Parameter-efficient Fine-tuning of MDGD via Gradient Masking
Parameter-efficient fine-tuning (PEFT) methods, such as adapters Houlsby et al. (2019) and LoRA Hu et al. (2021), aim to reduce the computational cost and memory usage when fine-tuning models on downstream tasks under practical constraints Han et al. (2024). However, due to the requirement of directly estimating gradient directions on the pre-trained model parameters, MDGD cannot be directly applied to these PEFT methods, which introduce additional model parameters whose gradients are separate from the original model weights.
To address this challenge, we propose a variant, MDGD-GM, by formulating the gradient regularization term in Eq. (8) as gradient masking that selects model weights with efficient gradient directions. Specifically, we define it as
$M_{\tilde{g}_{\theta}}=\mathbf{1}\left\{\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|\|\bar{g}_{\theta}\|}\geq T_{\alpha}\right\},$ (10)
@@ 3.3 Enabling Parameter-efficient Fine-tuning of MDGD via Gradient Masking
$\pi_{\theta}\leftarrow\pi_{\theta}-M_{\tilde{g}_{\theta}}\cdot\left(\nabla_{\theta}\mathcal{L}_{vl}(\theta)+\nabla_{\theta}\mathcal{L}_{v}(\theta)\right).$ (11)
@@
Table 1: Performance on various pre-trained tasks of LLaVA-1.5 models fine-tuned on Flickr30K and OKVQA. We report the best performance for each task in a bold font while the second best performance underlined.
4 Experiments
Datasets To evaluate catastrophic forgetting, we follow the setup of Zhu et al. (2024) and consider two models: LLaVA-1.5 (7B) and MiniCPM-V-2.0 (2.8B). For each model, we use a set of pre-trained tasks (including VQAv2, GQA, VizWiz, TextVQA, POPE, and MM-Bench) and fine-tuning tasks on previously unseen datasets (such as Flickr30k, OKVQA, TextCaps, and PathVQA). Full details on dataset composition are provided in the appendix.
Baselines We compare our approach against several baselines: standard fine-tuning following Zhu et al. (2024), LoRA-based fine-tuning (Hu et al., 2021), and Model Tailor (Zhu et al., 2024). For Model Tailor, we report the original results from their paper for comparison.
Implementation Details All experiments use the official Huggingface implementations of LLaVA-1.5 and MiniCPM-V-2.0, with LoRA adapters where applicable. Models are fine-tuned using BFloat16 precision on 2 NVIDIA A100 80GB GPUs. We include fine-grained implementation details in Appendix B.
LLM Usage In this paper, LLMs are only used for refining the writing of natural language.
4.1 Comparison Results
@@
LLaVA-1.5 adapts better to downstream tasks but is more prone to visual forgetting. We study the visual forgetting problem on the LLaVA-1.5 MLLM, and report performance comparison results in Table 1. We observe that the pre-trained LLaVA enables efficient instruction tuning on target tasks, where the zero-shot performance is near zero. When the model is fine-tuned on the image caption task, Flickr30K, which largely differs from the pre-trained tasks of visual question-answering, the model can learn a degraded multimodal representation, which causes visual forgetting in its projected visual representation space (in Section C). Fine-tuning on visual question-answering task OKVQA, which is similar to the pre-trained tasks, can also lead to MLLM’s visual understanding drift, due to the limited image-text pairs existing in the downstream task.
@@
Table 2: Performance on various pre-trained tasks of MiniCPM-V2.5 models fine-tuned on PathVQA and TextCaps. We report the best performance for each task in a bold font while the second best performance underlined.
MiniCPM-V-2.0 also experiences visual forgetting while limited in downstream task improvements. To validate the observation on a smaller MLLM, we report the comparison results of MiniCPM-V-2.0 with 2.8B model parameters in Table 2. We observe that compared with the LLaVA MLLM, MiniCPM suffers from less prominent visual forgetting. We attribute this observation to MiniCPM learning a more compact and constrained visual representation space during pre-training, causing the visual representations of target task images to be less aligned with those of the pre-trained MLLM. Consequently, MiniCPM exhibits limited improvement in downstream tasks, as its restricted ability to acquire additional visual knowledge leads to ineffective instruction tuning.
MDGD prevents visual forgetting while maintaining downstream task improvements. By employing MDGD in MLLM instruction tuning, we observe the LLaVA’s average performance drops on pre-trained tasks when fine-tuned on OKVQA and also improves when fine-tuned on Flickr30K, which demonstrates the efficiency of MDGD in mitigating visual forgetting. For the smaller MLLM, MiniCPM, MDGD achieves comparable fine-tuning improvements with direct fine-tuning, while completely eliminating visual forgetting in the pre-trained tasks. MDGD and its variants consistently achieve the best average performance for both MLLMs, demonstrating its great potential for incremental learning on individual downstream tasks.
Comparison with baseline methods. Table 1 shows that MDGD consistently outperforms both LoRA fine-tuning and Model Tailor Zhu et al. (2024) on LLaVA-1.5. LoRA suffers from visual forgetting due to projecting multimodal features into lower-rank spaces, especially on Flickr30K and OKVQA. Model Tailor, while effective for anti-forgetting in LLMs, is less robust for MLLMs and remains sensitive to the target dataset, performing better on Flickr30K than OKVQA. In contrast, MDGD achieves higher average scores and H-scores across datasets. In Table 2, MDGD improves average performance on MiniCPM tasks, reducing visual forgetting by 2.43% and 1.83% on PathVQA and TextCaps, respectively.
(a) $z^{vl}$ on PathVQA
(b) $\pi(X^{v})$ on PathVQA
(c) $z^{vl}$ on TextCaps
(d) $\pi(X^{v})$ on TextCaps
Figure 2: T-SNE plots of the distribution of extracted visual $\pi(X^{v})$ and multimodal $z^{vl}$ representations from pre-trained MiniCPM, and models with direct fine-tuning and MDGD on PathVQA and TextCaps.
Figure 3: Illustration of (a) the learning process of three methods based on task loss $\mathcal{L}_{vl}(\phi,\theta)$, (b) the average regularized cosine similarity $\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|\|\bar{g}_{\theta}\|}$ in Eq.(10) for gradient masking at varying ratios, and (c) the visual representation loss $\mathcal{L}_{v}(\phi,\theta)$ in Eq.(5) for gradient masking at varying ratios $\alpha$.
4.2 Ablation Study
Ablation study on visual alignment. We compare MDGD with its two variants, MDGD w/o visual align and MDGD-GM. MDGD w/o visual align enables MDGD without including visual representation loss $\mathcal{L}_{v}(\phi,\theta)$ Eq.(5), to understand the effect of directly optimizing to reduce the visual representation discrepancy between the current model and pre-trained model. We observe that MDGD w/o visual align maintains relatively comparable performance to MDGD on OKVQA and PathQA, due to the reduced need for visual representation adaptation in such visual question-answering tasks. In contrast, tasks like image captioning on Flickr30K and TextCaps benefit from feature alignment regularization, which directly mitigates visual understanding drift in the MLLM.
Ablation study on gradient masking. The other variant, MDGD-GM, leverages gradient masking to enable parameter-efficient fine-tuning (PEFT). We observe the PEFT variant of MDGD consistently achieves comparable performance across all tasks and backbone MLLMs, which only fine-tunes a subset of 10% original MLLM parameters used for direct fine-tuning and original MDGD. Different from conventional PEFT methods such as adapters, MDGD and its variants do not introduce additional parameters to the original model architecture, enabling incremental learning in an online setting (Maltoni and Lomonaco, 2019; Gao et al., 2023).
4.3 Representation Learning Analysis
T-SNE Analysis on Visual Representation To analyze the learning of visual and multimodal representation distributions in MLLMs, we create T-SNE Van der Maaten and Hinton (2008) plots to visualize the feature distributions extracted from pre-trained MLLMs, as well as MLLMs after standard fine-tuning and MDGD We illustrate the distributions of the multimodal features $z^{vl}$ extracted from the last token of the multimodal instruction tokens, and the visual features $\pi_{\theta}(X^{v})$ extracted from the last token of the input image tokens. We observe a consistent visual understanding drift in the MLLMs’ visual representation spaces after standard fine-tuning on PathVQA and TextCaps with MiniCPM (Figure 2b and 2d). By employing MDGD to mitigate visual forgetting, we observe that visual understanding drift is effectively reduced, allowing the fine-tuned MLLM to retain pre-trained visual capabilities.
We further observe a distributional discrepancy in the multimodal representation $z^{vl}$ of LLaVA (Figures 5a and 5c) between MDGD and the pre-trained MLLM. This discrepancy arises from the alignment of the MLLM to the target task through multimodal instructions, demonstrating effective adaptation to the downstream task of the LLaVA model. In addition, we also observe such multimodal distribution discrepancy reduces in a smaller MLLM, MiniCPM. This observation aligns with our findings on MiniCPM in Section 4.1, where we noted limited effects in model adaptation to downstream tasks. However, applying MDGD to MiniCPM mitigates visual forgetting by preventing degradation of both image and multimodal encodings into lower-rank representation spaces.
Effective Rank Analysis on Visual Representation To quantitatively analyze the visual forgetting problem (in Section C) described in Eq. (12), we calculate effective ranks of the visual representations extracted from the last hidden layer on the position of image tokens in individual MLLMs. We show the comparison results of LLaVA models in Figure 4(a) and MiniCPM models in Figure 4(b). We observe that with both the backbone models of LLaVA and MiniCPM, directly fine-tuning the pre-trained models on downstream tasks can lead to a consistent reduction of effective ranks in visual representations. Such observations validate the hypothesis in Section C regarding the potential visual forgetting problem in MLLM instruction tuning. In addition, we can observe that MDGD achieves consistent improvements in effective ranks compared with the standard fine-tuning method for both backbone MLLMs across various pre-trained tasks. In Figure 4(a), we observe that MDGD achieves comparable or even better effective ranks on pre-trained tasks, compared with the pre-trained LLaVA model. However, MDGD on MiniCPM in Figure 4(b) also suffers from the visual representation degradation problem, while MDGD consistently alleviates the problem. Such observation suggests a higher risk of visual forgetting in smaller-scale MLLMs.
(a) LLaVA models pretrained, finetuned, and fine-tuned with MDGD
(b) MiniCPM models pretrained, finetuned, and fine-tuned with MDGD
Figure 4: The effective rank comparison on individual downstream fine-tuning datasets.
4.4 Sensitivity Study
5 Conclusion
6 Limitation
In this work, we focus on MLLMs that process multimodal instructions consisting solely of visual and textual inputs. Given the limited availability of MLLMs across other modalities, our primary goal is to mitigate visual forgetting. However, our modality-decoupling approach is generalizable to other input modalities. Consistent with standard practices, we limit the instructions to two input modalities, though extending this to more diverse, free-form multimodal inputs remains an avenue for future research.
Acknowledgment
This work is partially supported by NSF IIS-2432486.
Appendix A T-SNE Analysis on LLaVA-1.5 Model
In addition to Section 4.3, we further include the T-SNE analysis on LLaVA-1.5 model. We observe a consistent visual understanding drift in the MLLMs’ visual representation spaces after standard fine-tuning on Flickr30K and OKVQA with LLaVA (Figure 5b and 5d).
@@ Acknowledgment
Figure 5: T-SNE plots of the distribution of extracted visual $\pi(X^{v})$ and multimodal $z^{vl}$ representations from pre-trained LLaVA-1.5, and models with direct fine-tuning and MDGD on OKVQA and Flickr30K.
Appendix B Implementation Details
  • Pre-trained Tasks: VQAv2 (Goyal et al., 2017), GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), SQA (Lu et al., 2022), TextVQA (Singh et al., 2019), POPE (Li et al., 2023c), and $M$-Bench (Liu et al., 2023).
  • Pre-trained Tasks: VizWiz, OKVQA, A-OKVQA Schwenk et al. (2022), Text-VQA, IconQA Lu et al. (2021), POPE, and $M$-Bench.
Baselines We compare our approach against several baselines:
  • Standard Fine-Tuning. For a fair comparison, we follow the setting of Model-Tailor Zhu et al. (2024), where LLaVA-1.5 is fine-tuned on the last 6 layers and its feature adapter, with a total of 1.2B parameters. MiniCPM is fine-tuned on the last 8 layers and its feature resampler, with 517M parameters.
Appendix C Visual Forgetting in MLLM Instruction-tuning
$Z\sim p(\cdot\mid X^{v},X^{l}),$
During instruction tuning, the model is optimized primarily to predict the target $y$. As described in Eq. (3), the IB objective introduces a trade-off between retaining visual information $I(X^{v};Z)$ and ensuring that $Z$ remains predictive of $y$ via $I(y;Z)$ Jiang et al. (2024a). In practice, however, instruction-tuning datasets are predominantly text-driven; thus, the learned visual representation $Z^{v}$ receives only indirect and often weaker supervision Wang et al. (2024).
$\mathcal{Z}_{\theta}^{v}=\left\{Z^{v}_{\theta}:Z\sim p_{\theta}(\cdot\mid X^{v},X^{l}),(X^{v},X^{l})\in D\right\},$
$\text{erank}(\mathcal{Z}_{\theta}^{v})<\text{erank}(\mathcal{Z}_{0}^{v}).$ (12)